OpenStudy (anonymous):

–2(x – 3) ≥ 5 – (x + 3) SOMEONE PLEASE EXPLAIN STEP BY STEP SMH...

OpenStudy (anonymous):

at the left side of that....use distribute multiplication while on the right side multiply -1 to x and 3 then do the operation

OpenStudy (anonymous):

–2(x – 3) ≥ 5 – (x + 3) okay so its look like that now -2x + 6 ≥ 5 -1x -3 like that?

OpenStudy (anonymous):

yes! you're correct! now combine like terms..

OpenStudy (anonymous):

how to do that?

OpenStudy (anonymous):

see those values with x? combine them at one side: either left or right...but I prefer it to be on the right side since -2x will be 2x which is greater than -x .. then those constant at the other.. NOTE: it is the same as –2(x – 3) = 5 – (x + 3)

OpenStudy (anonymous):

oh so we add -2x and -1x

OpenStudy (anonymous):

\[–2x + 6 ≥ 5 –x - 3\] \[ 6+3-5 \ge 2x - x\] \[4 \ge x\]

OpenStudy (anonymous):

how did you get that second part??

OpenStudy (anonymous):

notice that when you transfer a value to its' opposite side, it changes sign...

OpenStudy (anonymous):

\[–2x + 6 ≥ 5 –x - 3\] how did go from this to this \[ 6+3-5 \ge 2x - x\]

OpenStudy (anonymous):

like i said..simply combine those values with x and combine those values without x. now, i combined 6,3,and5. on the first part, notice that 3 is negative and when i transferred it to the left side it became positive... as when i transferred 5 to the left, notice that 5 became negative.

OpenStudy (anonymous):

oh kk i understand now what do we do??

OpenStudy (anonymous):

@unkabogable

OpenStudy (anonymous):

do the operation then it will be: \[4 \ge x\]

OpenStudy (anonymous):

thnx

OpenStudy (anonymous):

welcome! anytime! :)